# Chernoff bounds for branching random walks

### v4

@@ 2 Chernoff bound for BRW
In this paper, we study BRW from a new perspective. Traditionally, BRW is viewed as Galton-Watson tree weighted by spatial displacements and its study relies heavily on generating-function techniques. In contrast, we take BRW as RW ++ point process—maybe called “RWB”. In a random walk, viewed as a special BRW, produces exactly one child and is replaced by it. The intensity measure of the displacement point process $\{\xi_{1}\}$ is
@@
$\int_{\left|\,S_{n}-\mathbb{E}(S_{n})\,\right|\geq\lambda}\prod_{i=1}^{n}p_{i}(x_{i}\mid u_{i-1})\,dx_{1}dx_{2}\cdots dx_{n}=\Pr\left(|S_{n}-\mathbb{E}(S_{n})\geq\lambda|\right)$
@@
$\Pr\big(|Q_{n}(\alpha)-na|\geq\lambda\big)\leq 2e^{-\frac{1}{2}c\lambda^{2}/n}$
where $na$ is the expected position111nana” is used for “mean” in order to be consistent with the literature of BRW, not indicating a linear relationship with $n$ for an individual in nthnth generation, $Q_{n}(\alpha)$ is $\alpha$ quantile and
@@
Let $\tau$ be a tree of $Z$ leaves and $p_{(\tau)}(\lambda)$ be the proportion of positions which $\geq na+\lambda$ (in the $n$th generation), i.e.
$p_{(\tau)}(\lambda)=\frac{1}{Z}\sum_{i=1}^{Z}\mathbf{1}_{\{z_{i}(\tau)\geq na+\lambda\}}$
@@
$p_{(\tau)}(\lambda)\geq\alpha\iff Q_{n}(1-\alpha)\geq na+\lambda$ (2.5)
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=\Pr(S_{n}-na\geq\lambda)$
@@
$\mathbf{E}\left[p_{(\tau)}(\lambda)\right]=\frac{\sum_{i=1}^{Z}\mathbb{E}(\mathbf{1}_{\{z_{i}-na\geq\lambda\}})}{Z}=\frac{\sum_{1}^{Z}\Pr(S_{n}-na\geq\lambda)}{Z}=\Pr(S_{n}-na\geq\lambda)$ (2.6)
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=\Pr(S_{n}-na\geq\lambda)$
@@
$\mathbb{E}\left[\mathbf{1}_{\{z_{i}-na\geq\lambda\}}\right]$
is independent of the size of the tree because, by the hypothesis $(a)$, branching is independent of walking. Branching is also independent from leaf indexing so that $E(\mathbf{1}_{\{z_{i}-na\geq\lambda\}})$ is the same for any $i$. Below we will use $k$ instead of capital $Z$ for readability in summation.
@@
$\frac{1}{k}{\sum_{i=1}^{k}\mathbb{E}\left[\mathbf{1}_{\{z_{i}-na\geq\lambda\}}\right]}=\frac{1}{k}\sum_{i=1}^{k}\frac{1}{n_{k}}\sum_{\tau_{k}}\mathbf{1}_{\{z_{i}(\tau_{k})-na\geq\lambda\}}$ (2.7)
@@
$=\frac{1}{N_{t}}\sum_{k\geq 1}\sum_{\tau_{k}}\frac{\sum_{i=1}^{k}\mathbf{1}_{\{z_{i}(\tau_{k})-na\geq\lambda\}}}{k}$
$=\frac{1}{N_{t}}\sum_{\tau}\frac{\sum_{i=1}^{k}\mathbf{1}_{\{z_{i}(\tau)-na\geq\lambda\}}}{k}$
@@
$\alpha\cdot\Pr\Big(Q_{n}(1-\alpha)\geq na+\lambda\Big)\leq e^{-c\lambda^{2}/n}$
@@
$\Pr\Big(Q_{n}(1-e^{-\frac{1}{2}c\lambda^{2}/n})-na\geq\lambda\Big)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.8)
@@
$\Pr(Q_{n}(\alpha)-na\geq\lambda)\leq e^{-\frac{1}{2}c\lambda^{2}/n}\,;$ (2.9)
@@
$\Pr\Big(Q_{n}(\alpha)-na\leq-\lambda\Big)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.10)

### v5

@@ 2 Chernoff bound for BRW
In this paper, we study BRW from a new perspective. Traditionally, BRW is viewed as Galton-Watson tree weighted by spatial displacements and its study relies heavily on generating-function techniques. In contrast, we take BRW as RW ++ point process—maybe called “RWB”. In a random walk, viewed as a special BRW, a parent produces exactly one child and is replaced by it. The intensity measure of the displacement point process $\{\xi_{1}\}$ is
@@
$\int_{\left|\,S_{n}-\mathbb{E}(S_{n})\,\right|\geq\lambda}\prod_{i=1}^{n}p_{i}(x_{i}\mid u_{i-1})\,dx_{1}dx_{2}\cdots dx_{n}=\Pr\left(|S_{n}-\mathbb{E}(S_{n})|\geq\lambda\right)$
@@
$\Pr\big(|Q_{n}(\alpha)-n\mu|\geq\lambda\big)\leq 2e^{-\frac{1}{2}c\lambda^{2}/n}$
[…]
@@
Let $\tau$ be a tree of $Z$ leaves and $p_{(\tau)}(\lambda)$ be the proportion of positions which $\geq n\mu+\lambda$ (in the $n$th generation), i.e.
$p_{(\tau)}(\lambda)=\frac{1}{Z}\sum_{i=1}^{Z}\mathbf{1}_{\{z_{i}(\tau)\geq n\mu+\lambda\}}$
@@
$p_{(\tau)}(\lambda)\geq\alpha\iff Q_{n}(1-\alpha)\geq n\mu+\lambda$ (2.5)
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=\Pr(S_{n}-n\mu\geq\lambda)$
@@
$\mathbf{E}\left[p_{(\tau)}(\lambda)\right]=\frac{\sum_{i=1}^{Z}\mathbb{E}(\mathbf{1}_{\{z_{i}-n\mu\geq\lambda\}})}{Z}=\frac{\sum_{1}^{Z}\Pr(S_{n}-n\mu\geq\lambda)}{Z}=\Pr(S_{n}-n\mu\geq\lambda)$ (2.6)
@@
$\mathbb{E}\left[p_{(\tau)}(\lambda)\right]=\Pr(S_{n}-n\mu\geq\lambda)$
@@
$\mathbb{E}\left[\mathbf{1}_{\{z_{i}-n\mu\geq\lambda\}}\right]$
is independent of the size of the tree because, by the hypothesis $(a)$, branching is independent of walking. Branching is also independent from leaf indexing so that $E(\mathbf{1}_{\{z_{i}-n\mu\geq\lambda\}})$ is the same for any $i$. Below we will use $k$ instead of capital $Z$ for readability in summation.
@@
$\frac{1}{k}{\sum_{i=1}^{k}\mathbb{E}\left[\mathbf{1}_{\{z_{i}-n\mu\geq\lambda\}}\right]}=\frac{1}{k}\sum_{i=1}^{k}\frac{1}{n_{k}}\sum_{\tau_{k}}\mathbf{1}_{\{z_{i}(\tau_{k})-n\mu\geq\lambda\}}$ (2.7)
@@
$=\frac{1}{N_{t}}\sum_{k\geq 1}\sum_{\tau_{k}}\frac{\sum_{i=1}^{k}\mathbf{1}_{\{z_{i}(\tau_{k})-n\mu\geq\lambda\}}}{k}$
$=\frac{1}{N_{t}}\sum_{\tau}\frac{\sum_{i=1}^{k}\mathbf{1}_{\{z_{i}(\tau)-n\mu\geq\lambda\}}}{k}$
@@
$\alpha\cdot\Pr\Big(Q_{n}(1-\alpha)\geq n\mu+\lambda\Big)\leq e^{-c\lambda^{2}/n}$
@@
$\Pr\Big(Q_{n}(1-e^{-\frac{1}{2}c\lambda^{2}/n})-n\mu\geq\lambda\Big)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.8)
@@
$\Pr(Q_{n}(\alpha)-n\mu\geq\lambda)\leq e^{-\frac{1}{2}c\lambda^{2}/n}\,;$ (2.9)
@@
$\Pr\Big(Q_{n}(\alpha)-n\mu\leq-\lambda\Big)\leq e^{-\frac{1}{2}c\lambda^{2}/n}$ (2.10)